### v2

# Graph Neural Networks for O-RAN Mobility Management: A Link Prediction Approach

Abstract
Mobility performance has been a key focus in cellular networks up to 5G. To enhance Handover (HO) performance, 3GPP introduced Conditional Handover and Layer 1/Layer 2 Triggered Mobility (LTM) mechanisms in 5G. While these reactive HO strategies address the trade-off between HO failures and ping-pong effects, they often result in inefficient radio resource utilization due to additional HO preparations. To overcome these challenges, this article proposes a proactive HO framework for mobility management in O-RAN, leveraging user-cell link predictions to identify the optimal target cell for HO. We explore various categories of Graph Neural Networks (GNNs) for link prediction and analyze the complexity of applying them to the mobility management domain. Two GNN models are compared using a real-world dataset, with results demonstrating their ability to capture the dynamic and graph-structured nature of cellular networks. Finally, we present key insights from our study and outline future steps to enable the integration of GNN-based link prediction for mobility management in O-RAN networks.
Index Terms:
Graph Neural Networks, O-RAN, Handover, Mobility Management, Link Prediction.
I Introduction
Mobility management is a fundamental feature of cellular networks, ensuring seamless service continuity during user mobility by minimizing call drops, Radio Link Failures (RLFs) and unnecessary handovers (HOs), known as ping-ponging [1]. This management can be carried out through different HO mechanisms, which are responsible for decision-making and messages handling in the RAN. Traditionally, mobility management follows a reactive HO mechanism, where HO decisions rely on predefined network parameters. However, determining optimal mobility parameters remains complex due to the trade-off between HO failures and ping-pong effects. Additionally, the dynamic communication environment often limits reactive techniques. These challenges are further exacerbated in 5G networks, where densification, higher frequencies, and stringent URLLC requirements demand more sophisticated mobility management solutions [2].
The advent of Open Radio Access Network (O-RAN) has revolutionized the RAN by leveraging Artificial Intelligence (AI) or Machine Learning (ML) to enhance network performance and flexibility. These AI/ML models, deployed at the edge and centralized cloud, enable intelligent decision-making and automation in managing modern cellular networks. Mobility management is a key O-RAN use case, and recent studies have explored proactive HO mechanisms based on user trajectory [3], user measurements [4], and RLF prediction [5]. ML models such as sequence-to-sequence (Seq2Seq) and Long Short-Term Memory (LSTM) show promise in capturing long-term dependencies between user and cell metrics. However, these models process data sequentially, failing to leverage the graph-structured nature of communication networks, limiting generalization and reducing predictive performance on unseen networks and mobility patterns.
To address the challenges of modeling dynamic network topologies, Graph Neural Networks (GNNs) have become increasingly relevant in both academic and industrial context [6, 7]. GNNs operate on graph-structured data by iteratively aggregating information from neighboring nodes, making them suitable for tasks such as node or edge regression and link prediction. Prior work has applied GNNs to regression problems in mobility management. such as predicting signal quality or load-related metrics [8]. Although these outputs can support HO optimizations, the models stop at intermediate predictions, with the estimation of HO targets left as an unexplored work.
In this article, we study the capabilities of GNN link prediction for mobility management and how to overcome the technical challenges when adapting these ML models as proactive HO mechanism. This work is intended as an initial exploration, aimed at demonstrating the feasibility of using GNN-based link prediction to anticipate HOs in O-RAN systems. We propose next-cell prediction, a proactive HO mechanism that leverages a GNN-based link prediction model to forecast the upcoming cell connection for each user. We provide an in-depth analysis and model comparison of two categories of GNN link prediction models and analyze the complexity of applying them to the mobility management problem. Sec. II reviews the O-RAN framework, current HO mechanisms and the role of GNNs, while also summarizing recent advancements in proactive HO mechanisms. Sec. III outlines the problem formulation, proposes an O-RAN framework for next-cell prediction, and analysis in-depth the challenges of applying GNNs for mobility management. Sec. IV compares the performance of prediction accuracy of subgraph-based and autoencoder-based GNN link prediction models, trained on real-world datasets. Finally, Sec. V discusses the key findings and insights on the applicability of GNN-based link prediction for mobility management, concluding the article with practical implications and future directions.
II Background
II-A O-RAN
The O-RAN Alliance aims to create an intelligent, open, multi-vendor, interoperable, and autonomous radio access network. A core innovation of O-RAN is the introduction of Radio Intelligent Controllers (RICs). These include the non-Real-Time (non-RT) RIC for long-term optimization and planning, hosted within the Service Management and Orchestration (SMO). Complementing this is the near-Real-Time (near-RT) RIC for dynamic and real-time network configuration. O-RAN further fragments the 3GPP architecture, which already splits the Central Unit (CU) and the Distributed Unit (DU), by additionally dividing the CU into Control Plane (CU-CP) and User Plane (CU-UP).
II-B Handover mechanisms in O-RAN
Figure 1: Handover process in cellular networks.
This work addresses inter-gNB HO in O-RAN, utilizing the Xn interface for transitions between different gNBs, as shown in Fig. 1. The baseline HO process involves the User Equipment (UE) measuring neighboring cell metrics, such as Signal-to-Noise Ratio (SNR), and reporting these to its source gNB (1). If the source signal degrades and neighbor’s improves, the source gNB initiates the HO procedure. This involves a sequence of steps: an XnAP Handover Request is sent to the target gNB (2), the target gNB replies with an acknowledgment (3), the source gNB then instructs the UE to reconfigure its connection (4), a temporary user-plane tunnel is set up between source and target CU-UPs for data transfer, and finally the UE confirms the reconfiguration to the target gNB (6).
O-RAN facilitates the integration of advanced AI/ML-based solutions in the RICs via rApps and xApps. The distinction lies in their response time restrictions: rApps (non-RT RIC) operate with response times greater than 1 second, while xApps (near-RT RIC) require response times between 10 ms and 1 second. In the standard [9], O-RAN introduces the use of these AI/ML solutions to implement proactive HO mechanisms for mobility management. The non-RT RIC trains the AI/ML models using collected metrics via the O1 interface and then the deployed rApps sends A1 policies to the near-RT RIC. The xApp deployed in the near-RT RIC subsequently uses these A1 policies, alongside E2 measurements, to apply real-time configuration updates, facilitating proactive and adaptive HO decisions.
II-C GNNs for link prediction in mobility management
GNNs have emerged as powerful tools for learning over structured data. Their ability to process graph-structured data, by aggregating information from neighboring nodes, allows them to capture both local and global graph properties. Among the many tasks GNNs address (e.g., regression, classification), link prediction stands out as particularly relevant for scenarios where the goal is to infer potential or future connections between entities.
Several GNN architectures have been developed specifically for link prediction. Among them, encoder-decoder models such as Graph Autoencoders (GAE), which produce fixed embeddings. While Variational Graph Autoencoders (VGAE) model distributions, capturing uncertainty seen in dynamic mobility settings, subgraph-based methods focus on local neighborhoods around candidate links to capture short-range patterns, being particularly effective in sparse or evolving graphs. SEAL, for instance, applies GNNs to these subgraphs in a supervised manner, enabling fine-grained link prediction. Matrix factorization models like ComplEx embed nodes and relations into complex vector spaces, enabling them to capture asymmetry. Table I summarizes the most important GNN models for link prediction, outlining their inputs, outputs, benefits, and limitations.
TABLE I: Summary of link prediction models
Name Input Output Strengths Weaknesses
Autoencoders GAE Node features, edge features and edge indexes. Predicted existence or non-existence of edges. Its architecture is highly adaptable. Also, it is simpler and more efficient than other models. Unsupervised and deterministic model with a binary output, so no further analysis is allowed.
VGAE Node features, edge features and edge indexes. Predicted probabilities for each edge. Captures uncertainty and variability in the data. Improves the adaptability in networks with changing user behavior. Unsupervised model. Requires an additional process to determine an optimal threshold for the probabilities obtained.
Subgraph-based SEAL Enclosing subgraphs and node features. Binary classification (link or no link between two nodes). Focuses on local subgraphs, improving short-range predictions; supervised learning enhances performance. Subgraph extraction increases computational cost; requires labeled data.
ComplEx Node pairs and relation types (for multi-relational graphs). Predicted probabilities of links between nodes. Handles multi-relational graphs; captures asymmetric relationships effectively. Limited structural information beyond embeddings; computationally intensive for large-scale graphs.
II-D Existing solutions for mobility management
Recent AI/ML approaches for mobility management primarily use Seq2Seq models for optimizing intra-frequency HO based on UE trajectories [3], and LSTM models for incorporating mobility information into next-cell prediction [4]. While these methods show promising results, they rely on sequential data processing using Recurrent Neural Networks (RNN) or Reinforcement Learning (RL) approaches. These architectures are limited in their ability to capture the structural relationships present in cellular networks and often struggle to generalize across evolving topologies [10].
Cellular networks naturally form graphs of interdependent nodes with features, making GNNs suitable for solving network problems. They have been applied to radio resource management [2], 5G signal inference [11] and traffic prediction [6]. In the case of link prediction, models must account for factors like interference and load, which introduce temporal variability. Recent studies combine GNNs with RL, as in [8], where connection management is performed as a graph optimization task. While expressive, RL-based methods are often data-intensive and computationally expensive, limiting real-time applicability. Industry efforts such as Amazon’s GNN-based mobility solution [7] highlight the practical interest in this domain, though technical details remain unpublished. Furthermore, [5] combines GNNs and time-series transformers to predict RLFs, but focuses on link reliability than mobility management.
Despite these advancements, proactive HO management using GNNs remains underexplored. The complexity of dynamic graph-structured networks presents an opportunity for GNN-based methods adapted to mobility and operational constraints.
III Next-cell prediction for mobility management
III-A Problem formulation
The baseline HO mechanism can lead to signaling overhead and delays, due to necessary re-transmissions which cause the HO to be triggered too late. Additionally, the interruption time is typically between 50-90 ms, often too long for URLLC services. Network-triggered HO mechanisms like conditional HO and L1/L2 Triggered Mobility (LTM), although designed to improve this interruption time, can lead to ping-ponging. Moreover, triggers based on low-level metrics may ignore the overall network quality or congestion levels, potentially leading to suboptimal HO. Traditional ML proactive approaches, although promising, may struggle with adaptability and generalization in dynamic environments [10].
To address these limitations, we propose a GNN-based link prediction method for proactive mobility management in cellular networks. This approach distinguishes itself from traditional reactive or sequential methods by enabling direct next-cell prediction for UEs, thereby aiming to reduce signaling overhead and HO delays. By leveraging the spatial and graph-structured context of cellular networks, our models estimate future user-cell links. This initial study demonstrates the feasibility of supporting earlier HO decisions, with the potential to minimize ping-pong effects and improve HO timing. The autoencoder-based model, in particular, suggests a scalable solution for enhancing HO performance and user experience within O-RAN architectures.
III-B Next-Cell prediction rApp - A Walkthrough
Figure 2: Next-cell prediction framework for proactive HO.
The proposed O-RAN mobility management solution is shown on Fig. 2 and operates by integrating the next-cell prediction rApp into the Non-RT RIC. The workflow starts with UE collecting and transmitting measurement reports (1). Subsequently, both the serving and neighboring gNBs transmit RAN data and configurations, including L1/L2/L3 metrics, to the SMO Operations, Administration, and Maintenance (OAM) functions via the O1 interface (2). These OAM functions have a baseline measurement configuration to select the data of interest, and allows its access to the Non-RT RIC (3). A pre-trained GNN-based link prediction model, hosted as an rApp evaluates the collected RAN data to forecast the upcoming cell connections. Based on these predictions, the rApp generates A1 Policy update messages and transmits them to the Connection Management xApp via the A1 interface (4). This xApp, based on the the A1 policies and the E2 measurement metrics obtained from the E2 nodes (5.a and 5.b), generates RIC Control messages (6) to initiate radio resource re-allocation and HO request (7 and 8), mirroring the standard HO process from Fig. 1. Crucially, the only HO latency introduced by this solution is attributed to O1 data collection, model inference (evaluated in Sec. II-D), and policy generation, all of which are low-overhead operations. This design choice is fundamental to maintaining efficiency and adhering to the stringent timing requirements inherent in O-RAN deployments.
III-C Challenges of applying GNN models for mobility management
Applying GNN models for next-cell prediction in mobility management presents distinct challenges, primarily due to the dynamic, heterogeneous, and imbalanced characteristics of real-world cellular networks.
III-C1 Data heterogeneity
Cellular networks are heterogeneous, comprising diverse node types (e.g., gNBs, cells, UEs) and different edge types (e.g., radio links, HO transitions, control interfaces). This heterogeneity is critical for accurately modeling mobility patterns, as HO decisions depend on both user-cell associations and inter-cell relationships. However, standard GNN models for link prediction are designed for homogeneous graphs, which prevents the effective representation of these complex network structures.
III-C2 Data imbalance
Mobility management scenarios exhibit pronounced node imbalance, as the number of UE nodes significantly exceeds the number of cell nodes. This can lead to low connectivity and sparse graphs, which can significantly impact GNN models. The imbalance affects message-passing GNN models, since low-degree nodes may receive insufficient aggregated information, limiting predictive accuracy. For example, in next-cell prediction, imbalanced representations can bias the model toward frequent HO targets, reducing sensitivity to less common transitions. The imbalance also complicates prediction thresholds, as the models produce probabilistic scores rather than binary outputs. A default threshold of 0.5 is often unsuitable in imbalanced scenarios [12].
III-C3 Edge features
Edge features play a central role in modeling HO behavior, as they encode contextual information (e.g., signal strength, timing, HO frequency) critical to predicting next-cell transitions. However, many baseline GNN models aggregate node embeddings without considering edge attributes, limiting their effectiveness in O-RAN scenarios.
IV Model evaluation
This section evaluates two GNN models: VGAE based and SEAL. VGAE captures global embeddings, while SEAL focuses on localized subgraphs, providing a complementary comparison for mobility management. The following subsections detail the dataset, model creation, training methodology, and experimental setup.
IV-A Dataset description
The study utilizes an open-source dataset from Huawei Technologies [13], derived from a real-world operational network. It initially comprises approximately 35 million edges and 169,000 nodes (100,000 user nodes and 69,000 cell nodes). For privacy, all node and edge features are anonymized, represented as random properties normalized between 0 and 1, which limits the data analysis.
The data is structured across three main files:
  • node_cell: Contains 69,000 cell IDs, each with 1,000 associated feature columns
  • node_user: Includes 100,000 user IDs, each with 300 feature columns
  • edge_user_live_cell: Originally lists about 35 million edges, each specifying an edge ID, user ID, cell ID, and 37 edge-specific features. To manage potential redundancy and due to a lack of information on multiple edges, only one edge per user-cell pair (the one with the lowest ID) was used, reducing the total edges to approximately 7 million.
This dataset presents each user connected to 69 cells and each cell to 1,000 users, creating 1,000 isolated sub-networks without inter-connectivity. This structure yields an extremely sparse network density of 0.00049. Due to computational complexity, initial model development and feasibility studies were conducted on a subset of 985 cells and 10,000 users, with plans for gradual scaling. This specific dataset was selected for its scale, structure, and operational origin, providing a strong foundation for an initial GNN-based link prediction model development and feasibility study in O-RAN. Further validation across different datasets and network scenarios will be necessary to fully assess generalization performance.
IV-B Model creation
The link prediction task is formally defined over an attributed graph $G=(V,E,X),$ where $V$ denotes the set of nodes, $E$ the set of edges, and $X$ represents node and edge features. The objective is to predict the probability of a link existing between pairs of nodes $(u,v)$, corresponding to the likelihood that a user will handover to a given target cell.
Model adaptations were necessary due to the graph’s specific nature and the solution’s objectives. As the dataset comprises two node types connected through a single hop, the model design employs a single layer for data transmission between neighbors. Incorporating additional layers would result in redundant information transmission and lead to over-smoothing and learning inefficiencies. To address each challenges of GNN models for mobility management, the following adaptations are implemented:
1) Data heterogeneity: The heterogeneous topology is transformed into an equivalent homogeneous graph. It is carried out unifying UEs and cells in a single type of node, assigning a unique ID for each one while preserving relevant information, such as features and connections.
2) Data imbalance: Subgraph-based GNNs mitigate global data imbalance by capturing specific subgraph structural patterns. However, autoencoder-based models are particularly affected due to their reliance on neighboring embeddings, which can decrease overall link prediction accuracy. An attention-based mechanism, Graph Attention Network (GAT), is used to mitigate this by prioritizing important neighbors during aggregation, thereby improving representation for underrepresented nodes. Additionally, due to data imbalance, prediction thresholds in both models require empirical tuning to identify decision boundaries that balance true positives and true negatives, ensuring stable performance metrics (e.g., recall and accuracy).
3) Edge features: For subgraph-based models, edge features are incorporated at the initial stage of subgraph processing, being encoded with node features. GAT layers enable the inclusion of edge features in the autoencoder-based models.
Furthermore, specific adaptations are made for VGAE and SEAL models.
IV-B1 VGAE
The VGAE encoder’s message-passing operation, utilizing a single GAT layer (expressed as Eq. 1), where $\mathbf{H}^{(l-1)}$ is the previous layer node features, $\hat{\mathbf{A}}$ is the symmetrically normalized adjacency matrix, $\mathbf{H}^{(l)}$ is the output of the current layer, $\mathbf{W}^{(l)}$ is a trainable weight matrix, and $\sigma$ is a ReLu activation function. The GAT layer’s output is processed by two GCN layers to calculate the mean ($\mu$) and log standard deviation (logstd), capturing uncertainty in the graph structure. An inner product decoder then calculates edge probabilities, with a threshold applied for the classification of edges. This combination of VGAE and GAT focuses on influential nodes while adapting to dynamic changes, making it ideal for next-cell prediction.
$\mathbf{H}^{(l)}=\sigma\left(\hat{\mathbf{A}}\mathbf{H}^{(l-1)}\mathbf{W}^{(l)}\right),$ (1)
IV-B2 SEAL
In SEAL, the labeled subgraphs are processed by a GCN-based encoder, which learns a compact subgraph embedding. This enables the model to capture both structural and relational information from the dataset. The message-passing operation in SEAL can be expressed as Eq. 2, where $\mathbf{H}_{i}^{(l)}$ is the hidden representation of node $i$ in layer $l$, $\mathbf{W}^{(l)}$ is the trainable weight matrix at layer $l$, and $\sigma$ is a non-linear ReLu activation function.
$\mathbf{H}_{i}^{(l)}=\sigma\left(\sum_{j\in\mathcal{N}(i)}\mathbf{W}^{(l)}\mathbf{H}_{j}^{(l-1)}\right),$ (2)
IV-C Model training
A systematic training process was applied to optimize model performance. The dataset was divided into 80/10/10 training, validation, and test sets using a custom split function to ensure consistent, structure-preserving partitions. This split generates positive (existing) and negative (non-existing) edges for each set, which are necessary for contrastive supervision during training. This custom approach prevents potential issues associated with automatic split generation functions (e.g., PyTorch RandomLinkSplit), such as including existing edges as negative, and duplicated negative edges between different sets, which negatively impact model performance [14] [15].
VGAE and SEAL use labeled data for supervised training with Kullback-Leibler (KL) divergence and cross-entropy loss, respectively. The Adam optimizer was used for its efficiency in handling large and sparse datasets. Both models underwent fine-tuning on parameters such as weights, learning rate, latent representation size. During training, model performance was evaluated using Area Under the ROC Curve (AUC), Average Precision (AP), Precision, Recall, F1-score, Accuracy, and Matthews Correlation Coefficient (MCC). To address potential overfitting, early stopping was applied to both models, and L2 regularization in SEAL.
IV-D Performance comparison
Fig. 3 shows the mean metrics values obtained from 10 executions of each model. Both models, SEAL (blue) and VGAE (light orange), demonstrate comparable predictive performance over the 985 cells and 10,000 UEs graph. VGAE improves SEAL’s AP and Recall by 6.21% and 11.23%. Conversely, SEAL outperformed VGAE in AUC (4.5%), Precision (12.3%), F1-score (10.09%), Accuracy (0.61%), and MCC (2.81%). While VGAE effectively identifies edges within a broader context, it yields more False Positives (FP), thus excelling in FP-independent metrics like Recall. However, SEAL surpasses VGAE in other metrics due to VGAE’s higher FP count. SEAL’s use of localized subgraph information also enhanced its precision in predictions with smaller datasets.
The two darker orange bars in Fig. 3 examine VGAE scalability by increasing the graph size to 10,000 cells and 20,000 cells, both with the 100,000 UEs. SEAL could not run on these larger datasets due to the significant computational cost associated with subgraph creation. VGAE performance improved overall with the increase of the dataset size, from VGAE-985 to VGAE-10k. Recall showed highest growth (4.91%), followed by AUC (3%), Accuracy (2.2%) and F1-score (2.18%). This reflects an increase in correctly predicted edges, due to improved generalization from training the model on a broader variety of subgraph structures and interactions. As a result, the model learns to assign more accurate attention to nodes during training. However, VGAE-20k model performance deteriorates. This occurred because the dataset maintained the same number of UEs, but duplicated cells, leading to a drastic increase in node density, which likely affects the attention mechanism and increases FPs (key limitation of the model).
Fig. 4 shows the models training times. VGAE outperforms SEAL in execution speed, achieving a 98.84% reduction in training time with the 985 cells dataset due to its direct graph processing. This computational burden prevented SEAL from executing on the two large datasets. VGAE also demonstrates faster convergence during training, achieving satisfactory results within less than 50 epochs. This rapid convergence, coupled with its simpler architecture, enables efficient training, even on larger datasets. Furthermore, VGAE’s training time increases linearly with the graph size as it performs embeddings for each node of the network.
Fig.5 illustrates VGAEs reduction in SEAL’s inference time with the 985 cells dataset, also attributed to its direct execution without subgraphs. VGAEs inference time, similar to its training time, also grows linearly with dataset size. Crucially, all model inference times are below or close to 1 second, demonstrating their high efficiency and suitability for O-RAN rApp deployments, which are characterized by operational response times greater than 1 second.
Figure 3: Performance metrics obtained with VGAE and SEAL, along with different dataset sizes.
Figure 4: Training time of VGAE and SEAL, along with different dataset sizes.
Figure 5: Inference time of VGAE and SEAL, along with different dataset sizes..
The findings demonstrate the efficacy of graph-based methodologies for link prediction in large-scale networks, even without features or temporal information, relying solely on graph structure for learning. While prior studies used GNN-based regressions models for HO estimation (e.g., [5] for radio link failure and [3] for user mobility), no prior work directly applied GNN models for link prediction in this context. Although not directly comparable, our proposal achieved an F1-score 5% higher (84%) than [5] (79%), using only graph structure information. Similarly, [3] reported 85% accuracy with nodes features processing, improvable to 90% by including user’s historical mobility data. This suggests that including temporality and features would enhance our model performance, which already demonstrates high efficiency using only the graph structure for training.
V Discussion and limitations of this work
This section discusses the main results of applying GNN-based link prediction for mobility management in O-RAN, along with the limitations of our work.
V-A GNN-link prediction and trade-offs
The choice between SEAL and VGAE involves critical trade-offs for O-RAN mobility management. While SEAL is more selective in edge identification, reducing FPs crucial for minimizing unnecessary HOs and system costs, its high computational complexity limits scalability. In contrast, VGAE is computationally more efficient and better suited for large-scale networks due to its attention mechanism, exhibiting significantly faster training and inference times. However, VGAE’s performance can decline in highly dynamic environments or with increased node density, as shown in the 20k cell dataset scenario.
Real-world wireless networks are dynamic, with new UEs and shifting data distributions. As both models are transductive, they might struggle to generalize to unseen UEs and out-of-distribution inputs, highlighting the need for inductive capabilities. These challenges highlight the need for future research into inductive Spatio-Temporal GNNs, such as TGNs or GraphMixer, which can incorporate dynamic user movement patterns and network dynamics to enhance accuracy and generalization. Including temporal information in datasets is essential to fully leverage these advanced models for more precise and timely HO decisions in real-word O-RAN deployments.
V-B Lack of available datasets
A significant limitation of this work is its reliance on a single real-world dataset, whose anonymized features and structure constrain data analysis and limit the generalization of our findings. Access to large-scale, realistic datasets with temporal information remains a critical barrier for O-RAN mobility management research. Existing publicly available datasets, such as those from testbeds (e.g., Colosseum) or simulations (e.g., ns-O-RAN), lack the necessary scale, diversity, or temporal resolution for comprehensive mobility modeling and GNN training, often featuring too few UEs or relying on approximated models. Future work will focus on training and evaluating models on datasets with temporal user mobility traces, potentially leveraging emulation tools like Keysight EXata or RICTest to benchmark against other HO solutions.
V-C Low connectivity and node density
Low connectivity and node density are inherent challenges in cellular network, where users are typically connected to a single serving cell. This sparsity significantly impacts GNN models, particularly VGAE and GAE, by hindering accurate link prediction due to insufficient neighboring information. The issue is compounded by the isolated subgraphs observed in our dataset, where users connect to individual cells without broader inter-cell connectivity. A potential solution involves incorporating higher-level nodes, such as CU or DU nodes, to increase graph density, interconnect cells, and facilitate richer message passing, thereby improving embedding generation.
V-D Negative edges creation
The process of creating negative edges is critical for autoencoder-based models, as highlighted by [14, 15]. Relying on random sampling functions can erroneously introduce positive edges as negative, or duplicate negative edges across datasets, leading to incorrect training or biased evaluation metrics. Specifically, including test edges in the training set can cause overfitting and distribution shift. To mitigate these issues, future work could leverage the bipartite nature of cellular networks during negative edge creation, which may enhance the attention mechanism for cell nodes and reduce VGAE performance degradation in large-scale environments.
VI Conclusions
In this article we proposed a GNN-based link prediction solution for proactive mobility management in O-RAN. Using real-world cellular data, we compared two GNN models: an autoencoder-based model and a subgraph-based model where results show the autoencoder model offers superior speed and training efficiency, especially for larger datasets, achieving competitive predictive performance with high recall and average precision. While the subgraph model excelled in precision, the autoencoder model demonstrated better generalization and recall with larger datasets, though its performance decreased with high node density. Given the ability of the autoencoder-based model to deliver accurate next-cell predictions efficiently and its potential to address challenges in mobility management, we advocate for further exploration of this approach. Future research could refine and extend these findings, paving the way for designing innovative GNN-based HO mechanisms in future networks, marking a significant paradigm shift in mobility management.
Acknowledgment
This project is partially funded by the European Union-NextGenerationEU within the Framework of the Project ”Massive AI for the Open RadIo b5G/6G Network (MAORI)”.
Biographies
Ana González Bermúdez (M.Sc.’2022): Researcher at Laude.
Miquel Farreras (M.Sc.’2019, Ph.D.’2024): Researcher at Laude.
Milan Groshev (M.Sc.’2016, Ph.D.’2022): Senior Researcher at Laude.
José Antonio Trujillo (M.Sc.’2020): Ph.D Candidate at UMA.
Isabel de la Bandera (M.Sc.’2009, Ph.D.’2017): Associate professor at UMA.
Raquel Barco (M.Sc.’1997, Ph.D.’2007): Full professor at UMA.

### v3

@@
 We explore various categories of Graph Neural Networks (GNNs) for link prediction and analyze the complexity of applying them to the mobility management domain.
-Two GNN models are compared using a real-world dataset, with results demonstrating their ability to capture the dynamic and graph-structured nature of cellular networks.
-Finally, we present key insights from our study and outline future steps to enable the integration of GNN-based link prediction for mobility management in O-RAN networks.
-Index Terms:
-Graph Neural Networks, O-RAN, Handover, Mobility Management, Link Prediction.
+Two GNN models are compared using an anonymized real-world dataset and a synthetic dataset obtained with Keysight’s EXata emulation tool.
+The results show the models’ ability to capture the dynamic and graph-structured nature of cellular networks.
+Finally, we present key considerations for real-world deployment that outline future steps to enable the integration of GNN-based link prediction for mobility management in O-RAN networks.
 I Introduction
 Mobility management is a fundamental feature of cellular networks, ensuring seamless service continuity during user mobility by minimizing call drops, Radio Link Failures (RLFs) and unnecessary handovers (HOs), known as ping-ponging [1].
-This management can be carried out through different HO mechanisms, which are responsible for decision-making and messages handling in the RAN.
+This management can be carried out through different HO mechanisms, which are responsible for decision-making and message handling in the Radio Access Network (RAN).
 Traditionally, mobility management follows a reactive HO mechanism, where HO decisions rely on predefined network parameters.
@@
 These AI/ML models, deployed at the edge and centralized cloud, enable intelligent decision-making and automation in managing modern cellular networks.
-Mobility management is a key O-RAN use case, and recent studies have explored proactive HO mechanisms based on user trajectory [3], user measurements [4], and RLF prediction [5].
-ML models such as sequence-to-sequence (Seq2Seq) and Long Short-Term Memory (LSTM) show promise in capturing long-term dependencies between user and cell metrics.
+Mobility management is a key O-RAN use case, and recent studies have explored proactive HO mechanisms based on user trajectory [3] and User Equipment (UE) measurements [4].
+ML models such as sequence-to-sequence (Seq2Seq) and Long Short-Term Memory (LSTM) show promise in capturing long-term dependencies between UE and cell metrics.
 However, these models process data sequentially, failing to leverage the graph-structured nature of communication networks, limiting generalization and reducing predictive performance on unseen networks and mobility patterns.
-To address the challenges of modeling dynamic network topologies, Graph Neural Networks (GNNs) have become increasingly relevant in both academic and industrial context [6, 7].
+To address the challenges of modeling dynamic network topologies, Graph Neural Networks (GNNs) have become increasingly relevant in both academic and industrial contexts [5, 6].
 GNNs operate on graph-structured data by iteratively aggregating information from neighboring nodes, making them suitable for tasks such as node or edge regression and link prediction.
-Prior work has applied GNNs to regression problems in mobility management.
-such as predicting signal quality or load-related metrics [8].
-Although these outputs can support HO optimizations, the models stop at intermediate predictions, with the estimation of HO targets left as an unexplored work.
+Prior work has applied GNNs to regression problems in mobility management, such as predicting signal quality or load-related metrics [7].
+Although these outputs can support HO optimizations, the models stop at intermediate predictions, with the estimation of HO targets remaining unexplored.
 In this article, we study the capabilities of GNN link prediction for mobility management and how to overcome the technical challenges when adapting these ML models as proactive HO mechanism.
@@
 Sec.
-III outlines the problem formulation, proposes an O-RAN framework for next-cell prediction, and analysis in-depth the challenges of applying GNNs for mobility management.
+III outlines the problem formulation, proposes an O-RAN framework for next-cell prediction, and analyzes in-depth the challenges of applying GNNs for mobility management.
 Sec.
-IV compares the performance of prediction accuracy of subgraph-based and autoencoder-based GNN link prediction models, trained on real-world datasets.
+IV compares the performance of prediction accuracy of subgraph-based and autoencoder-based GNN link prediction models, trained on real-world and synthetic datasets.
+Sec.
+V summarizes the main considerations for the real-world deployment of GNN-based link prediction for mobility management.
 Finally, Sec.
-V discusses the key findings and insights on the applicability of GNN-based link prediction for mobility management, concluding the article with practical implications and future directions.
+VI concludes the article with key findings and future directions.
 II Background
 II-A O-RAN
-The O-RAN Alliance aims to create an intelligent, open, multi-vendor, interoperable, and autonomous radio access network.
-A core innovation of O-RAN is the introduction of Radio Intelligent Controllers (RICs).
-These include the non-Real-Time (non-RT) RIC for long-term optimization and planning, hosted within the Service Management and Orchestration (SMO).
-Complementing this is the near-Real-Time (near-RT) RIC for dynamic and real-time network configuration.
-O-RAN further fragments the 3GPP architecture, which already splits the Central Unit (CU) and the Distributed Unit (DU), by additionally dividing the CU into Control Plane (CU-CP) and User Plane (CU-UP).
+O-RAN is an industry effort to open and disaggregate the RAN, enabling multi-vendor interoperability, software driven control, and AI/ML assisted automation.
+Its architecture centers on two RAN Intelligent Controllers (RICs), the Non-Real-Time (Non-RT) RIC hosted in the Service Management and Orchestration (SMO) system, and the Near-Real-Time (Near-RT) RIC close to the RAN.
+rApps run in the Non-RT to train models and issue policies over the A1 interface, while xApps run in the Near-RT RIC to apply those policies together with measurements received over the E2 interface.
+Network configuration and performance data flow between the SMO and RAN through the O1 interface.
+In parallel, functional disaggregation separates the Central Unit (CU) and the Distributed Unit (DU) and further splits the CU into control plane (CU-CP) and user plane (CU-UP) parts, which simplifies scaling, upgrades, and vendor mixing within a gNB deployment.
 II-B Handover mechanisms in O-RAN
@@
 1.
-The baseline HO process involves the User Equipment (UE) measuring neighboring cell metrics, such as Signal-to-Noise Ratio (SNR), and reporting these to its source gNB (1).
+The baseline HO process involves the UE measuring neighboring cell metrics, such as Signal-to-Noise Ratio (SNR), and reporting these to its source gNB (1).
 If the source signal degrades and neighbor’s improves, the source gNB initiates the HO procedure.
@@
 O-RAN facilitates the integration of advanced AI/ML-based solutions in the RICs via rApps and xApps.
-The distinction lies in their response time restrictions: rApps (non-RT RIC) operate with response times greater than 1 second, while xApps (near-RT RIC) require response times between 10 ms and 1 second.
+The distinction lies in their response time restrictions: rApps operate with response times greater than 1 second, while xApps require response times between 10 ms and 1 second.
 In the standard  [8], O-RAN introduces the use of these AI/ML solutions to implement proactive HO mechanisms for mobility management.
-The non-RT RIC trains the AI/ML models using collected metrics via the O1 interface and then the deployed rApps sends A1 policies to the near-RT RIC.
+The Non-RT RIC trains the AI/ML models using collected metrics via the O1 interface and then the deployed rApp sends A1 policies to the Rear-RT RIC.
 The xApp deployed in the near-RT RIC subsequently uses these A1 policies, alongside E2 measurements, to apply real-time configuration updates, facilitating proactive and adaptive HO decisions.
@@
 II-D Existing solutions for mobility management
-Recent AI/ML approaches for mobility management primarily use Seq2Seq models for optimizing intra-frequency HO based on UE trajectories [3], and LSTM models for incorporating mobility information into next-cell prediction [4].
-While these methods show promising results, they rely on sequential data processing using Recurrent Neural Networks (RNN) or Reinforcement Learning (RL) approaches.
-These architectures are limited in their ability to capture the structural relationships present in cellular networks and often struggle to generalize across evolving topologies [10].
+Recent AI/ML approaches for mobility management [9] include Seq2Seq models for intra-frequency HO optimizations using UE trajectories [3], LSTM models for next-cell prediction with mobility context [4] and Deep Reinforcement Learning (DRL) methods that learn policies from interactions [10, 9].
+These families of models capture temporal patterns and can adapt to changing traffic and load.
+In this article, we study a graph-based formulation that operates on the network structure, with the goal of assessing whether link prediction can support proactive decisions in an O-RAN.
+We propose this as complementary to DRL and recurrent models, as each family targets different aspects of the problem.
 Cellular networks naturally form graphs of interdependent nodes with features, making GNNs suitable for solving network problems.
@@
 Recent studies combine GNNs with RL, as in [7], where connection management is performed as a graph optimization task.
-While expressive, RL-based methods are often data-intensive and computationally expensive, limiting real-time applicability.
 Industry efforts such as Amazon’s GNN-based mobility solution [6] highlight the practical interest in this domain, though technical details remain unpublished.
@@
 These OAM functions have a baseline measurement configuration to select the data of interest, and allows its access to the Non-RT RIC (3).
-A pre-trained GNN-based link prediction model, hosted as an rApp evaluates the collected RAN data to forecast the upcoming cell connections.
+A pre-trained (with prior knowledge of the network) GNN-based link prediction model, hosted as an rApp evaluates the collected RAN data to forecast the upcoming cell connections.
 Based on these predictions, the rApp generates A1 Policy update messages and transmits them to the Connection Management xApp via the A1 interface (4).
@@
 Crucially, the only HO latency introduced by this solution is attributed to O1 data collection, model inference (evaluated in Sec.
-II-D), and policy generation, all of which are low-overhead operations.
+IV.D), and policy generation, all of which are low-overhead operations.
 This design choice is fundamental to maintaining efficiency and adhering to the stringent timing requirements inherent in O-RAN deployments.
@@
 VGAE captures global embeddings, while SEAL focuses on localized subgraphs, providing a complementary comparison for mobility management.
-The following subsections detail the dataset, model creation, training methodology, and experimental setup.
-IV-A Dataset description
-The study utilizes an open-source dataset from Huawei Technologies [13], derived from a real-world operational network.
-It initially comprises approximately 35 million edges and 169,000 nodes (100,000 user nodes and 69,000 cell nodes).
-For privacy, all node and edge features are anonymized, represented as random properties normalized between 0 and 1, which limits the data analysis.
-The data is structured across three main files:
-  • node_cell: Contains 69,000 cell IDs, each with 1,000 associated feature columns
-  • node_user: Includes 100,000 user IDs, each with 300 feature columns
-  • edge_user_live_cell: Originally lists about 35 million edges, each specifying an edge ID, user ID, cell ID, and 37 edge-specific features.
-To manage potential redundancy and due to a lack of information on multiple edges, only one edge per user-cell pair (the one with the lowest ID) was used, reducing the total edges to approximately 7 million.
-This dataset presents each user connected to 69 cells and each cell to 1,000 users, creating 1,000 isolated sub-networks without inter-connectivity.
-This structure yields an extremely sparse network density of 0.00049.
-Due to computational complexity, initial model development and feasibility studies were conducted on a subset of 985 cells and 10,000 users, with plans for gradual scaling.
-This specific dataset was selected for its scale, structure, and operational origin, providing a strong foundation for an initial GNN-based link prediction model development and feasibility study in O-RAN.
-Further validation across different datasets and network scenarios will be necessary to fully assess generalization performance.
+The following subsections detail the , model creation, training methodology, and experimental setup.
+IV-A Used datasets
+To validate the model performance we used two datasets: i) an open-source Real-World dataset (RW dataset) from Huawei Technologies [^1] https://snap.stanford.edu/data/telecom-graph.html, and ii) a synthetic EMulated dataset (EM dataset) generated using Keysight 5G network emulator EXata [^2] https://www.keysight.com/fr/en/product/SN100VGLA/exata-network-modeling-5g.html.
+RW was selected for its scale, structure, and operational origin, providing a strong foundation for an initial GNN-based link prediction model development.
+The small-scale EXata dataset (EM), was generated for validation of the models generalization capabilities across diverse network scenario with real radio metrics.
+The datasets main characteristics are:
+  • RW is composed of 35 million edges and 169,000 nodes (100,000 UEs and 69,000 cells), which allows the analysis of the models’ scalability.
+All nodes and edge have features associated: UEs have 1,000, cells have 300, and edges have 37.
+However, for privacy reasons, all node and edge features are anonymized, represented as random properties normalized between 0 and 1, which limits the data analysis.
+  • EM is composed of 449,152 edges and 101 nodes (70 UEs and 31 cells).
+The emulator allows the collection of L1/L2 UE radio metrics.
+Performing feature engineering on the collected data and excluding the highly correlated features, as UE-cell edges we used: RSRP, RSRQ, Transport Blocks, Available Resources Blocks, Packet Size, MCS, Signal Power Uplink, and Signal Power Downlink.
+From the raw data, we create the same graph data structure for both datasets.
+Due to computational complexity of processing RW and the scale differences, initial development is conducted on a subset of 985 cells and 10,000 users, with further for gradual scaling.
+Also, RW limits its interpretability due to a lack of information on multiple edges between the same pair of UE-cell.
+To manage potential redundancy, only one edge per user-cell pair (the one with the lowest ID) was used, reducing the total edges to approximately 7 million.
+Same procedure is then used in EM, reducing the edges to 489.
+This way, RW has each UE connected to 69 cells and each cell to 1,000 UEs, creating 1,000 isolated sub-networks without inter-connectivity.
+EM has each UE connected with different number of cells, ranging from two to six.
+These structures yields sparse network densities of 0.00049 and 0.09683, respectively.
 IV-B Model creation
 The link prediction task is formally defined over an attributed graph $G=(V,E,X),$ where $V$ denotes the set of nodes, $E$ the set of edges, and $X$ represents node and edge features.
-The objective is to predict the probability of a link existing between pairs of nodes $(u,v)$, corresponding to the likelihood that a user will handover to a given target cell.
+The objective is to predict the probability of a link existing between pairs of nodes $(u,v)$, corresponding to the likelihood that a user will HO to a given target cell.
 Model adaptations were necessary due to the graph’s specific nature and the solution’s objectives.
@@
 Incorporating additional layers would result in redundant information transmission and lead to over-smoothing and learning inefficiencies.
-To address each challenges of GNN models for mobility management, the following adaptations are implemented:
+To address each challenge of GNN models for mobility management, the following adaptations are implemented:
 1) Data heterogeneity: The heterogeneous topology is transformed into an equivalent homogeneous graph.
@@
 A systematic training process was applied to optimize model performance.
-The dataset was divided into 80/10/10 training, validation, and test sets using a custom split function to ensure consistent, structure-preserving partitions.
+The datasets was divided into 80/10/10 training, validation, and test sets using a custom split function to ensure consistent, structure-preserving partitions.
 This split generates positive (existing) and negative (non-existing) edges for each set, which are necessary for contrastive supervision during training.
-This custom approach prevents potential issues associated with automatic split generation functions (e.g., PyTorch RandomLinkSplit), such as including existing edges as negative, and duplicated negative edges between different sets, which negatively impact model performance [14] [15].
+This custom approach prevents potential issues associated with automatic split generation functions (e.g., PyTorch RandomLinkSplit), such as including existing edges as negative, and duplicated negative edges between different sets, which negatively impact model performance [15].
 VGAE and SEAL use labeled data for supervised training with Kullback-Leibler (KL) divergence and cross-entropy loss, respectively.
@@
 Fig.
-3 shows the mean metrics values obtained from 10 executions of each model.
-Both models, SEAL (blue) and VGAE (light orange), demonstrate comparable predictive performance over the 985 cells and 10,000 UEs graph.
-VGAE improves SEAL’s AP and Recall by 6.21% and 11.23%.
-Conversely, SEAL outperformed VGAE in AUC (4.5%), Precision (12.3%), F1-score (10.09%), Accuracy (0.61%), and MCC (2.81%).
-While VGAE effectively identifies edges within a broader context, it yields more False Positives (FP), thus excelling in FP-independent metrics like Recall.
-However, SEAL surpasses VGAE in other metrics due to VGAE’s higher FP count.
-SEAL’s use of localized subgraph information also enhanced its precision in predictions with smaller datasets.
-The two darker orange bars in Fig.
-3 examine VGAE scalability by increasing the graph size to 10,000 cells and 20,000 cells, both with the 100,000 UEs.
+3 shows the mean metrics values obtained from 10 executions of each model-dataset combination.
+The blue bars correspond to the results with the EM dataset with SEAL (SEAL-EM) and VGAE (VGAE-EM).
+The orange bars represent SEAL and VGAE applied to the smallest partition of RW, with 985 cells and 10,000 UEs (SEAL-985-RW and VGAE-985-RW), and the VGAE results when scaling the dataset to 10,000 cells and 100,000 UEs (VGAE-10k-RW) and to 20,000 cells and 1,000,000 UEs (VGAE-20k-RW).
+Using the EM dataset, VGAE-EM outperforms SEAL-EM in AUC (0.81%), AP(18.63%), Precision (6.08%), F1-score (0.1%) and Accuracy (1.36%), whereas SEAL-EM outperforms VGAE-EM in Recall (7.53%).
+Meanwhile, using the 985-RW dataset, VGAE-985-RW achieves higher AP (6.21%) and Recall (11.23%) than SEAL-985-RW, but SEAL-985-RW outperforms VGAE-985-RW in AUC (4.5%), Precision (12.3%), F1-score (10.09%) and Accuracy (0.61%).
+The differing outcomes observed in the models’ performance, depending on the used dataset, arise from their inherent differences in density and size.
+SEAL shows lower performance with the EM, as the low dataset size produces small subgraphs that limit its learning capacity.
+In contrast, VGAE enhances its performance with EM, as it is almost 200 times denser than RW, which particular benefits to autoencoder-based models capturing edges within a broader context.
+This is the reason for VGAE worsening with RW, generating more False Positives (FP) due to the sparsity of the graph.
+However, with the larger dataset SEAL achieves stronger performance in FP-sensitive metrics.
+Fig.
+3 also illustrates VGAE scalability with increasing graph size.
 SEAL could not run on these larger datasets due to the significant computational cost associated with subgraph creation.
-VGAE performance improved overall with the increase of the dataset size, from VGAE-985 to VGAE-10k.
-Recall showed highest growth (4.91%), followed by AUC (3%), Accuracy (2.2%) and F1-score (2.18%).
+VGAE performance improved overall when scaling from VGAE-985-RW to VGAE-10k-RW, with Recall showing the highest growth (4.91%), followed by AUC (3%), Accuracy (2.2%) and F1-score (2.18%).
 This reflects an increase in correctly predicted edges, due to improved generalization from training the model on a broader variety of subgraph structures and interactions.
 As a result, the model learns to assign more accurate attention to nodes during training.
-However, VGAE-20k model performance deteriorates.
+However, VGAE-20k-RW model performance deteriorates.
 This occurred because the dataset maintained the same number of UEs, but duplicated cells, leading to a drastic increase in node density, which likely affects the attention mechanism and increases FPs (key limitation of the model).
@@
 4 shows the models training times.
-VGAE outperforms SEAL in execution speed, achieving a 98.84% reduction in training time with the 985 cells dataset due to its direct graph processing.
+VGAE outperforms SEAL in execution speed, achieving a 98.84% reduction in training time for 985-RW and a 38.5% for EM due to its direct graph processing.
 This computational burden prevented SEAL from executing on the two large datasets.
-VGAE also demonstrates faster convergence during training, achieving satisfactory results within less than 50 epochs.
+VGAE also demonstrates faster convergence during training, achieving satisfactory results within less than 50 epochs in both datasets.
 This rapid convergence, coupled with its simpler architecture, enables efficient training, even on larger datasets.
 Furthermore, VGAE’s training time increases linearly with the graph size as it performs embeddings for each node of the network.
-Fig.5 illustrates VGAEs reduction in SEAL’s inference time with the 985 cells dataset, also attributed to its direct execution without subgraphs.
+Fig.5 illustrates VGAEs reduction in SEAL’s inference time for 985-RW and for EM, also attributed to its direct execution without subgraphs.
 VGAEs inference time, similar to its training time, also grows linearly with dataset size.
 Crucially, all model inference times are below or close to 1 second, demonstrating their high efficiency and suitability for O-RAN rApp deployments, which are characterized by operational response times greater than 1 second.
-Figure 3: Performance metrics obtained with VGAE and SEAL, along with different dataset sizes.
-Figure 4: Training time of VGAE and SEAL, along with different dataset sizes.
-Figure 5: Inference time of VGAE and SEAL, along with different dataset sizes..
-The findings demonstrate the efficacy of graph-based methodologies for link prediction in large-scale networks, even without features or temporal information, relying solely on graph structure for learning.
+Figure 3: Performance metrics obtained with VGAE and SEAL, along with different dataset types and sizes.
+Figure 4: Training time of VGAE and SEAL, along with different dataset types and sizes.
+Figure 5: Inference time of VGAE and SEAL, along with different dataset types and sizes.
+The findings demonstrate the efficacy of graph-based methodologies for link prediction in different networks, even without feature information (RW dataset) or temporal processing, relying solely on graph structure for learning.
 While prior studies used GNN-based regressions models for HO estimation (e.g.,  [12] for radio link failure and [3] for user mobility), no prior work directly applied GNN models for link prediction in this context.
@@
 This suggests that including temporality and features would enhance our model performance, which already demonstrates high efficiency using only the graph structure for training.
-V Discussion and limitations of this work
-This section discusses the main results of applying GNN-based link prediction for mobility management in O-RAN, along with the limitations of our work.
-V-A GNN-link prediction and trade-offs
-The choice between SEAL and VGAE involves critical trade-offs for O-RAN mobility management.
-While SEAL is more selective in edge identification, reducing FPs crucial for minimizing unnecessary HOs and system costs, its high computational complexity limits scalability.
-In contrast, VGAE is computationally more efficient and better suited for large-scale networks due to its attention mechanism, exhibiting significantly faster training and inference times.
-However, VGAE’s performance can decline in highly dynamic environments or with increased node density, as shown in the 20k cell dataset scenario.
-Real-world wireless networks are dynamic, with new UEs and shifting data distributions.
-As both models are transductive, they might struggle to generalize to unseen UEs and out-of-distribution inputs, highlighting the need for inductive capabilities.
-These challenges highlight the need for future research into inductive Spatio-Temporal GNNs, such as TGNs or GraphMixer, which can incorporate dynamic user movement patterns and network dynamics to enhance accuracy and generalization.
-Including temporal information in datasets is essential to fully leverage these advanced models for more precise and timely HO decisions in real-word O-RAN deployments.
-V-B Lack of available datasets
-A significant limitation of this work is its reliance on a single real-world dataset, whose anonymized features and structure constrain data analysis and limit the generalization of our findings.
-Access to large-scale, realistic datasets with temporal information remains a critical barrier for O-RAN mobility management research.
-Existing publicly available datasets, such as those from testbeds (e.g., Colosseum) or simulations (e.g., ns-O-RAN), lack the necessary scale, diversity, or temporal resolution for comprehensive mobility modeling and GNN training, often featuring too few UEs or relying on approximated models.
-Future work will focus on training and evaluating models on datasets with temporal user mobility traces, potentially leveraging emulation tools like Keysight EXata or RICTest to benchmark against other HO solutions.
-V-C Low connectivity and node density
-Low connectivity and node density are inherent challenges in cellular network, where users are typically connected to a single serving cell.
-This sparsity significantly impacts GNN models, particularly VGAE and GAE, by hindering accurate link prediction due to insufficient neighboring information.
-The issue is compounded by the isolated subgraphs observed in our dataset, where users connect to individual cells without broader inter-cell connectivity.
-A potential solution involves incorporating higher-level nodes, such as CU or DU nodes, to increase graph density, interconnect cells, and facilitate richer message passing, thereby improving embedding generation.
-V-D Negative edges creation
-The process of creating negative edges is critical for autoencoder-based models, as highlighted by [14, 15].
-Relying on random sampling functions can erroneously introduce positive edges as negative, or duplicate negative edges across datasets, leading to incorrect training or biased evaluation metrics.
-Specifically, including test edges in the training set can cause overfitting and distribution shift.
-To mitigate these issues, future work could leverage the bipartite nature of cellular networks during negative edge creation, which may enhance the attention mechanism for cell nodes and reduce VGAE performance degradation in large-scale environments.
+V Considerations for real-world deployment
+Having demonstrated the feasibility of using GNNs for HO prediction, deploying this solution in a live O-RAN environment requires addressing several key challenges.
+The following considerations are crucial for evolving the proposed framework from a proof-of-concept into a robust, scalable, and adaptable real-world tool.
+V-A Enhancing Model Architecture for Dynamic Networks
+A primary consideration for real-world deployment is the evolution of the model to handle the dynamic and temporal nature of cellular networks.
+The VGAE and SEAL models are transductive.
+They are trained on a fixed graph and struggle to generalize to new UEs entering the network or to different network environments without retraining.
+Furthermore, a key challenge is modeling the continuous evolution of a UE’s connection as it moves, which constantly changes the state of its wireless channel to the serving cell.
+This can be represented in our framework by creating multiple, time-stamped edges between the same UE-cell pair.
+Thereby, the model can be fed with the data related with the edge evolution over the time, and predict the edge to a new cell.
+Future work should explore the adaptation of inductive models, such as Spatio-Temporal GNNs, which learn transferable patterns applicable to unseen nodes.
+This type of models can natively learn from the sequence of the channel changes to better understand a UE’s trajectory and make more context-aware predictions.
+V-B Addressing Data-Centric Challenges
+The unique characteristics of cellular network data introduce further challenges that must be managed for real-world deployment.
+Cellular graphs are naturally sparse, with UEs connected to the serving cell.
+This low graph connectivity can hinder the message-passing mechanism essential for GNNs, as nodes have limited information to aggregate from their neighbors.
+A potential solution is to enrich the graph structure by incorporating higher-level network nodes, such as CUs and DUs, to create a more densely connected graph that facilitates richer information flow.
+Another critical data-centric task is the generation of negative edges for training, as a naive random sampling approach can introduce flaws like including existing links as negative examples or duplicating samples, leading to overfitting and biased evaluation.
+Future implementations should use a structure-aware sampling strategy to generate more realistic negative samples.
+Finally, GNNs demonstrate that they can operate effectively on anonymized data by learning from the graph structure.
+However, GNNs present a ”black box” when it comes down to the model explainability.
+This limitation can be addressed by integrating specialized explainability GNN frameworks, such as GNNExplainer or GraphSHAP designed to interpret GNN predictions.
+V-C Ensuring Scalability and Operational Viability
+For integration of our solution into a real-world scenario, scalability is the central requirement.
+The solution must remain computationally efficient, operate within the available budget, and grow with network size.
+GNN architectures differ in cost and accuracy.
+Subgraph based models such as SEAL can be precise on small datasets, yet their per sample search and training cost rises quickly, which limits scale.
+Models such as VGAE are lighter and fit better for large graphs, although performance can drop in very dense areas.
+To meet operator needs, a modular approach can be used.
+The operator network can be partitioned into regions, for example dense urban zones with frequent HOs, suburban zones with moderate mobility, and high speed corridors with fast mobility.
+Each region runs a model and set of hyperparameters tailored to its conditions, which keeps inference latency and memory within target limits.
 VI Conclusions
 In this article we proposed a GNN-based link prediction solution for proactive mobility management in O-RAN.
-Using real-world cellular data, we compared two GNN models: an autoencoder-based model and a subgraph-based model where results show the autoencoder model offers superior speed and training efficiency, especially for larger datasets, achieving competitive predictive performance with high recall and average precision.
+Using real-world and synthetic cellular data, we compared two GNN models: an autoencoder-based model and a subgraph-based model, with results showing that the autoencoder model offers superior speed and training efficiency, especially for larger datasets, achieving competitive predictive performance with high recall and average precision.
 While the subgraph model excelled in precision, the autoencoder model demonstrated better generalization and recall with larger datasets, though its performance decreased with high node density.